At maximum speed an airplane travels 2,460 miles against the wind in 6 hours. Flying with the wind the plane can travel the same distance in 5 hours. What is the speed of the plane with no wind?
Answer by mananth(16946)

let speed of wind be y mph
let speed of plane in still air be x
..
Against wind the speed = x-y
distance traveled = 2460
2460/x-y= 6
6x-6y= 2460
x-y=410...........1
..
with wind speed = x+y
distance is same 2460
2460/x+y=5
5x+5y=2460
x+y=492............2
..
Add equation 1 & 2
x-y+x+y=410+492
2x=902
x= 451 mph speed of plane in still air.
..
Plug the value of x in equation 1
451-y=410
y=451-410
y=40 mph the speed of wind
.
Your feed back about the solution is very important in improvement.